OpenStudy (anonymous):

Which fraction expresses –30/54 in lowest terms? A. – 10/27 B. – 10/18 C. –15/27 D. – 5/9

OpenStudy (micahm):

openstudy.com/updates/4fe57619e4b06e92b873b448

OpenStudy (catlover5925):

D

OpenStudy (anonymous):

I believe its gonna be d.

OpenStudy (anonymous):

Whenever I reduce fractions I try to half them first, and keep halfing them until it can't be reduced anymore. So here we go:\[-\frac{ 30 }{ 54 }\] divide by two on top and bottom\[-\frac{ 15 }{ 27 }\] Now we can't half them anymore, so I'm going to divide the numerator and denominator by 3, because they can still be reduced further\[-\frac{ 5 }{ 9 }\] This can't be reduced any further. So D is the correct answer.